A COUPLED-CHANNEL CALCULATION OF THE B^{3}Σ_{u}^{−} − X^{3}Σ_{g}^{−} SPECTRUM OF S_{2}

\underline{S. T. GIBSON}, S. J. CAVANAGH, and B. R. LEWIS, Research School of Physical Sciences and Engineering, The Australian National University, Canberra ACT 0200, Australia (email: Stephen.Gibson@anu.edu.au); G. STARK, Physics Department, Wellesley College, Wellesley, MA 02481.

Diatomic sulfur (S_{2}) has been observed in absorption, via the strong ultraviolet B^{3}Σ_{u}^{−} − X^{3}Σ_{g}^{−} system, in the atmospheres of Jupiter and Io. However, the spectrum is not well understood. We apply a coupled-channel Schrödinger equation (CSE) model developed for the VUV B^{3}Σ_{u}^{−}−X^{3}Σ_{g}^{−} transition of O_{2}^{a} to the analogous S_{2} spectrum, employing the potential curves and coupling scheme of Wheeler et al.,^{b} as a starting point. The CSE calculations provide valuable insight into the photodissociation spectrum, defining the roles of the key predissociating electronic states and demonstrating complex interference effects that are more pronounced than for O_{2}.

aB. R. Lewis, S. T. Gibson, F. T. Hawes, and L. W. Torop, Phys. Chem. Earth C 26, 519 (2001).

bM. D. Wheeler, S. M. Newman, and A. J. Orr-Ewing, J. Chem. Phys. 108, 6594 (1998).